OpenStudy (anonymous):

If the money and interest are left in the bank, how much is the total accumulated savings at the end of the second year for a savings of $3,500 at 8% annually compounded interest? A. $3,742.30 B. $3,895.53 C. $3,935.10 D. $4,082.40

OpenStudy (tkhunny):

What have you tried?

OpenStudy (anonymous):

I did the depreciation formula but that was definitely wrong. I'm not exactly sure how to solve this.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

It's a savings account. It should grow. 3500*(1.08)^2 =

OpenStudy (anonymous):

Where did you get the 1.08? @tkhunny

OpenStudy (tkhunny):

The problem statement. "8% annually compounded interest"

OpenStudy (anonymous):

@tkhunny Oh, i thought it would be 0.08, not 1.08

OpenStudy (tkhunny):

If you wish to calculate Simple Interest and use I = p*r*t, then you would use t = 0.08. This is compound interest and we're no longer in that world. The Compound Accumlation factor associated with 8% interest compounded annually, is 1.08.

OpenStudy (anonymous):

Ohhhh ok, thanks so much @tkhunny!!

OpenStudy (anonymous):

@tkhunny so the answer is D?

OpenStudy (tkhunny):

That's what I get. Do you doubt?

OpenStudy (anonymous):

@tkhunny Nope! Cause that's the answer I got! But I have a question...

OpenStudy (anonymous):

Is this the formula I'm supposed to use for this kind of problem? A= P(1+r)^t

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Yes. That is a single deposit (P), accumulated at some periodic interest rate (r) for a specified number of periods, (t). In your problem, we have P = 3500, r = 0.08, and t = 2

OpenStudy (anonymous):

Oh okay! Thanks again @tkhunny :)